OpenStudy (lisa123):

Find all the rational zeros of the function. x^3-31x-30

OpenStudy (dayakar):

let p(x) =x^3-31x -30 find p(-1)

OpenStudy (lisa123):

0

OpenStudy (lisa123):

p(x)=0

OpenStudy (dayakar):

right (x-1) is a factor of p(x)

OpenStudy (dayakar):

divide p(x) with (x-1)

OpenStudy (lisa123):

how did you know to use -1?

OpenStudy (dayakar):

sum of the coefficients odd powers x = sum of the coefficients even powers of x then x-1 is a factor

OpenStudy (dayakar):

is it clear

OpenStudy (dayakar):

what happens

OpenStudy (lisa123):

can you use the rational roots theorem?

OpenStudy (dayakar):

is it remainder theorem

OpenStudy (lisa123):

no I mean to find all the possible roots

OpenStudy (dayakar):

ok , can u divide given expression with (x-1)

OpenStudy (dayakar):

sorry divide with x+1

OpenStudy (lisa123):

yes

OpenStudy (dayakar):

write the answer

OpenStudy (lisa123):

x^2-x-30

OpenStudy (dayakar):

very good

OpenStudy (dayakar):

now split middle term x^2-6x+5x-30

OpenStudy (dayakar):

x(x-6)+5(x-6) =(x-6)(x+5) these are other two factors

OpenStudy (lisa123):

so all of the factors are -1,6,-5, 0 ?

OpenStudy (dayakar):

p(x)= x^3-31x-30 =(x+1)(x^2-x-30) +(x+1)(x-6)(x+5)

OpenStudy (dayakar):

x+1,(x-6),(x+5) are factors

OpenStudy (dayakar):

-1, 6 and -5 are roots

OpenStudy (dayakar):

is it clear

OpenStudy (lisa123):

Yea thank you @dayakar